OpenStudy (anonymous):

A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet. a. Find the length of the side of the lot opposite the 60° angle.

OpenStudy (anonymous):

Well, \[\sin \theta=\frac{ opposite }{ hypotaneuse }\] Since you know what sin(30) is and what the opposite equals, you can use that to find the length of the hypotaneuse. Then you can use \[\cos \theta=\frac{ adjacent }{ hypotaneuse }\] where you know Cos(30) and the hypotaneuse. This will give you the value you seek.

OpenStudy (anonymous):

yeah I got that but I am coming up with 82

OpenStudy (anonymous):

I'm getting 82 for the hypotaneuse.

OpenStudy (anonymous):

okay, then I must be on the right track just it is getting a little jumbled

OpenStudy (anonymous):

so would the sine be o.5

OpenStudy (anonymous):

Mhm, sin(30)=1/2

OpenStudy (anonymous):

no no 0.5000

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I'm sorry, I'm not sure what you mean. You're asking if sin(30)=.5?

OpenStudy (anonymous):

a. The side opposite the 60° is 71.0141 feet b. The hypotenuse is 82 feet c. sin(30)=0.5000 cos(30)=0.8660 tan(30)=0.5773

OpenStudy (anonymous):

Yes, that's right.

OpenStudy (anonymous):

just wasn't sure

OpenStudy (anonymous):

thank you